OpenStudy (anonymous):

Find the surface area of a square pyramid that has a base of 2 feet and a slant height of 2.5.

OpenStudy (anonymous):

I think I did this correct and the answer in the lesson is wrong. Here's my calculation. Surface Area=LA+B Lateral Area=1/2Perimeter*slant height LA=1/2 (8)*(2.5) LA=4*2.5 LA=10 Surface Area=LA+B Surface Area=10+8 Surface area=18